OpenStudy (anonymous):

At a musical, 400 tickets were sold. Adult tickets sold for $8 and student tickets sold for $5. If the total amount collected was $2450, how many adult tickets were sold?

OpenStudy (anonymous):

a+s=400 8a+s*5=2450

OpenStudy (anonymous):

solve for a

OpenStudy (anonymous):

Irene do you know how to solve this?

OpenStudy (anonymous):

No, that's why i'm asking how to solve it. I don't know how to put it into an equation.

OpenStudy (anonymous):

You can solve using a)substitution b)elimination

OpenStudy (anonymous):

For this problem, sub is easier

OpenStudy (anonymous):

Okay, I did that. I solved for variable a and got a decimal.

OpenStudy (anonymous):

let's use elimination instead a+s=400 8a+s*5=2450 -5a-5s=-2000 8a+5s=2450

OpenStudy (anonymous):

I multiplied the first equation by -5 so variable 's' cancel out

OpenStudy (anonymous):

add two equation together -5a-5s=-2000 8a+5s=2450 ------------------ 3a=450 a=450/3

OpenStudy (anonymous):

Any question on what I did?

OpenStudy (anonymous):

How did ou get -2000?

OpenStudy (anonymous):

*you

OpenStudy (anonymous):

I multiplied this equation ( a+s=400 ) by -5

OpenStudy (anonymous):

Solved using substitution(another method( 1)a+s=400 2)8a+5s=2450 Rearrange the first equation s=400-a plug value of s into second equation 8a+5s=2450 8a+5(400-a)=2450 8a+2000-5a=2450 3a=450 a=450/3

OpenStudy (anonymous):

Any question?

OpenStudy (anonymous):

How do you get -5 in the first place?

OpenStudy (anonymous):

I chose to multiply by -5 so that when I add the two equation, variable 's' would cancel

OpenStudy (anonymous):

You can multiply an equation by any number as long as you do it on both side of equality

OpenStudy (anonymous):

Ohhhhh! Got it.